OpenStudy (anonymous):

Amit wants to estimate 1 1/9 + 3 6/7. To estimate, he simplifies the expression 1+4. Is Amit’s estimate a close estimate, an underestimate, or an overestimate? A. underestimate B. overestimate C. close estimate

OpenStudy (anonymous):

@CREEPYMLPMASTER

OpenStudy (anonymous):

@zBrandz23

OpenStudy (anonymous):

@jojo4eva

OpenStudy (anonymous):

@CaptainBatSparrow

OpenStudy (anonymous):

@Evil-cow-8

OpenStudy (anonymous):

@Ankh

OpenStudy (anonymous):

@AnimeAsgardian

OpenStudy (anonymous):

You'll need to convert those mixed fractions to improper fractions and solve it. Do you know how to convert them?

OpenStudy (anonymous):

wait let me do it

OpenStudy (anonymous):

If you converted it, you don't have to solve it till the end.

OpenStudy (anonymous):

i cant do it @itsbribro

OpenStudy (itsbribro):

im sorry i dont know it :(

OpenStudy (anonymous):

1 1/9 = 10/9 = 1.1 3 6/7 = 27/7 = 3.8 1.1+3.8 = 4.9 C. Close Estimate